A camera & case costs $310.00. The Camera costs $300 more than the case and the case costs the price of the set minus the cost of the camera. you want to buy the case only. how much change would you get back from $100
Found 2 solutions by briannadee, JoelSchwartz:
Answer by briannadee(2)

The camera and the case cost $310.00 total.
The camera by its self cost $300.00.
It says the case costs the price of the set minus the cost of the camera. So it means that the case costs $310.00 - (the cost of the camera) $300.00 which would equal $10.00.
The case costs $10.00.
And you would get back $90.00 in change out of $100.

Answer by JoelSchwartz(130)

c=cost of camera
x=cost of case
c=300+x
The cost of the camera represented by c is equal to 300 dollars plus the cost of the case which is represented as x.
x=310-c
The cost of the case represented as x equals the cost of the set minus the cost of the camera. The cost of the set is 310.
x=310-(300+x)
x=310-300-x
x=10-x
2x=10
x=5
Because the cost of the case is five dollars you subtract five from one hundred and get 95 dollars back.